Data set: ABP=90/60 mm Hg, HR 110 beats/min, cardiac output = 3 L/min, CVP = 15 mm Hg, PAP = 50/35 mm Hg, PCWP = 6 mm Hg, EF = 60%.

Please show all work to justify answer

1. Calculate the PVR
Answer #1

Pulmonary vascular resistance:

PVR= 80 ( MPAP- PCWP)/ CO

MPAP = Mean pulmonary artery pressure.

PWAP = Central venous pressure that is pulmonary capillary wedge pressure.

CO= Cardiac Output.

MPAP= Systolic pulmonary B.P +2( Diastolic pulmonary B.P)/3

= 50+ 2× 35/3

= 50+ 70/3

= 120/3

MPAP = 40mm.hg

Apply in PVR calculation,

= 80( MAP-PCWP)/CO

= 80( 40- 6) /3

= 80[34]/3

= 80×34/3

= 2720/3

= 906.66

Answer= PVR= 906.66